an irregular variation is one which occurs predictably and at random, usually over a very short time period. it is impossible to allow for irregular variations in forecasting rather, they are used to explain abnormalities after the event.
hope this helps :-)
Expected number of heads is 1/4 * 32 or 8 heads.
What is the probability of getting a red king?
Assuming you're drawing from a complete shuffled 54-card deck, chances are one in twenty-seven.
2 dice are rolled find the probability of getting exactly 2?
Each die can come up in 6 ways.
You get exactly 2 only if both dice come up as a ' 1 '.
The probability of the first one coming up '1' is (1/6).
The probability of the second one coming up as a '1' is (1/6).
The probability of both coming up as '1' is (1/6 x 1/6) = (1/36) = 2.78% (rounded)
If you toss 3 coins what are the odds of getting exactly 3 heads?
The odds of getting 3 heads in a toss of 3 coins is 1 in 8, or 0.125.
Each coin is probabalistically unrelated to each other, so you simply multiply the odds for each coin. 1 in 2 times 1 in 2 times 1 in 2 is 1 in 8.
If you roll two dice what are the odds that you get a one or a five?
I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.
I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.
I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.
I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.
i THINK IT IS .05
Every discipline of mathematics has contemporary problems and theories being worked on. Some examples of the more popular mathematics topics being worked on now are Lie algebras, knot theory, and game theory.
Find the chance of getting more than 15 in one throw with 3 dice?
How many outcomes are there if you spin three six sided number cubes?
216... there are 6 outcomes for each cube, so 6^3 is 216
When rolling two six sided dice what is the probability of rolling a one and a four?
There are 36 possible outcomes (6 X 6). the probability of getting a one is 1/6; the probability of getting a four is 1/6. 1/6 X 1/6 = 1/36. Then, you multiply 1/36 by 2 because you can either get a 4 and a 1 or a 1 and a 4. Therefore, the answer is:
2/36 = 1/18
Statistics what is the probality that on five rolls of a fair die you roll three or more 1s?
Since there are 6 numbers (1-6) on each die, and you're rolling it 5 times, the total number of possibilities is 65. You now have to find out, out of the total, how many possibilities fill your requirements. For 5 rolls of one the chance is 1/65. There are 5 possible sets of rolls where you'll get 4 1s (4 1s and a 2, 3, 4, 5, or 6). Lastly there are 25 possible sets of rolls where you'll get 3 1s (3 1s and a 2 and 2 or a 2 and 3 or a 2 and 4 etc). so adding up all the possibilities that is 31/65. That is a percentage of .4% or 4 out of 1000 rolls of all the dice you will get 3, 4, or 5 1s. Sounds like a yahtzee question.
What are the odds of rolling an even number on a die?
4-6 Chances
-----------------------------------------------------------
From Rafaelrz.
There are three even numbers in a regular die, 2, 4 and 6.
There are six equally probable outcomes of which three are
unfavourable.
So the odds of rolling an even number are 3:3 or 1:1.
This means there is a 50% probability of success.
What is the z score for a given raw score?
Let your raw score be x and M the mean and S the standard deviation.
The Z score for your specific x is Z=(x-M)/S
So say your score is 80 (out of 100) and the mean is 70 and the standard deviation is 10.
Then the z score for your 80 is:
(80-70)/10=1
If on the other hand you got a 60, then the z score would be -1.
As n increases the sampling distribution of pˆ (p hat) becomes approximately normal.
How did dice get numbers on them?
It happened after dice had been used in game playing for almost 3,000 years. Sometime around
that epoch, or close to it, the dice manufacturers, at their annual convention, had to face the
situation that they had all ignored and repressed for all those centuries ... the game was just
not popular any more, and their sales of dice ... never a market leader ... were now totally sinking
through the floor. I mean they were all taking the pipe. Where originally there had been a dice
manufacturer for every two or three caves, now there were whole cities without a single one.
In spite of repeated government stimulus, bailout, welfare, and designation as historic structures,
these factories had steadily closed in droves throughout the length and breadth of recorded
human history, and they continued to do so. It wasn't only a matter of financial survival either ...
if you could see the kind of mindless drivel that had replaced the game of dice, you'd understand
that it was just as much a matter of honor, of manhood, of self-esteem. Wherever you looked,
whole families were turning away from their dice, getting more excitement and tittilation out of
burying stones, chewing their toenails, and throwing apples in the river. A pall hung over
that year's dice manufacturers' convention, like a pall. They knew that someone had to do
something in a hurry, or this convention would be their last; next year at convention time,
they'd have to spend the week at the office, or ... even worse ... stay home.
It was against this backdrop of seeming hopelessness that Murray ... a newcomer to the industry ...
mounted his campaign for the presidency of the dice manufacturers' trade association, there at the
annual convention. It didn't take much for Murray to run away with the election. He had the vote
locked up the first time he took the podium, and pointed out to everyone that even though there was
still plenty of crap, they weren't even doing it with dice at their own convention ! Murray's solution
to the whole problem was a simple one, as some of the best and most revolutionary ideas in history
have often been. Instead of the classic six-sided, symmetrical, featureless, polished cubes that
they had been manufacturing and peddling for centuries, Murray explained how a total, cataclysmic
turnaround in the dice market would follow a simple modification to the product ... assigning a
number from 1 to 6 to each face, and carving the corresponding number of shallow depressions
into the face.
My tale has already exceeded any amount of the dear reader's time and attention that I could
reasonably have requested, and so it must end without further details. It remains only to mention
that Murray obviously had it right. Dice became what they is today. The game, the game piece, and
the production lines that continue to bring it to us, were all revived. And if there remain any who
would doubt the value of Murray's contribution to the modern game, they need only consider how it
might be different if dice still had no spots. See what I mean ?
58=390625
There are five choices for each of the questions. To find the total number of ways to answer multiply 5 for each of the eight questions.
What is the standard error if the population standard deviation is 100 and the sample size is 25?
Formula for standard error (SEM) is standard deviation divided by the square root of the sample size, or s/sqrt(n). SEM = 100/sqrt25 = 100/5 = 20.
What is the probability a girl is chosen if 15 students are girls and 10 students are boys?
The ratio of girls to total students is 15:25, or 3:5. Three out of five students are girls so there would be a 60% probability that a girl would be chosen; a 2 out of 5 chance, or 40% probability that a boy would be chosen.
How many sums are there with two dice?
If you roll two dice and add the numbers together, you might get 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, so there are eleven possibilities total. One is not included in this list because that would mean the first dice would roll a one and the second dice would roll a...zero? That's impossible!
7/128, or about 5.5%
The student has a 1/2 probability of getting each question correct. The probability that he passes is the probability that he gets 10 correct+probability that he gets 9 correct+probability that he gets 8 correct:
P(passes)=P(10 right)+P(9 right)+P(8 right)=[(1/2)^10]+[(1/2)^10]*10+[(1/2)^10]*Combinations(10,2)=[(1/2)^10](1+10+45)=56/1024=7/128.
What is a possible outcome for a protostar that does no have enough material?
A brown dwarf.
A brown dwarf.
A brown dwarf.
A brown dwarf.